40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 7 15:46:42 2013, Jenne, Update, Locking, PR2 flipped 
    Reply  Fri Feb 8 02:12:14 2013, Jenne, Update, Locking, PRMI work 
       Reply  Fri Feb 8 02:38:04 2013, Koji, Update, Locking, PRMI work 
       Reply  Fri Feb 8 11:01:18 2013, Jamie, Update, Locking, PRMI work 
          Reply  Fri Feb 8 12:39:32 2013, yuta, Update, Locking, PRMI work 
          Reply  Fri Feb 8 15:53:48 2013, Jenne, Update, Locking, PRMI work 
             Reply  Fri Feb 8 17:41:34 2013, Jenne, Update, Locking, PRMI work 
                Reply  Fri Feb 8 19:39:02 2013, Koji, Update, Locking, PRMI work 
                Reply  Fri Feb 8 20:05:15 2013, Jenne, Update, Locking, PRMI work 
Message ID: 8030     Entry time: Fri Feb 8 02:12:14 2013     In reply to: 8024     Reply to this: 8031   8032
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PRMI work 

[Jenne, Yuta]

Lots of work, no solid conclusions yet.  In-vac, we aligned MICH and the PRM.  Out of vac, we got beam on AS and REFL paths.  We can lock MICH, but we're not as happy with PRCL.

 

In-vac alignment: 

To get the beam centered on TT2 in yaw, Koji helped us out and moved TT1 with the sliders a little bit. Then to get the beam centered on PRM and PR2, Koji moved the TT2 sliders a little bit.

Yuta and I then moved PR2 forward a few mm, to keep the optical path length of the PRC approximately (within ~1mm, hopefully) the same as always.  After my PR2 optic swapping earlier, the pitch alignment was no longer good.  I loosened the screws holding the wire clamp to the optic holder, and tapped it back and forth until the alignment was good.  Of course, the screw-tightening / pitch-checking is a stochastic process, but eventually we got it.  A small amount of yaw adjustment by twisting the PR2 TT was also done, but not much was needed.

Beam was a little off in pitch at ITMY, so Yuta poked the top of PR3, and that one single poke was perfect, and the beam was very nicely centered on the ITMY target.  Beam was getting through BS target just fine.  We checked at ITMX, and the beam looked pretty centered, although we didn't put in a target.  We didn't do anything to BS while we were in-vac, since it was already good.

We aligned the ITMs so that their beams were retroreflecting back to the BS.  After this, we saw nice MICH fringes.

We aligned the PRM so that its beam was retroreflecting.

We checked that we were getting REFL and AS beams out of the vacuum, which we were (a small amount of adjustment was done to AS path steering mirrors).

AS table alignment:

We did a bit of tweaking of the REFL path, and lots of small stuff to the AS path. 

The AS beam was coming out of vac at a slightly different place in yaw, so we moved the first out of vac AS steering mirror so the beam hit the center, rather than ~1/3 of the way to the edge.  We then aligned the beam through the lens, to the camera, and to AS55.  Most significantly, we removed the BS that was just before AS55.  This was sending beam to a dump, but it is in place to send beam over to AS110, once we get back to real locking.  We measured ~30 microwatts of power going to the AS55 PD, while MICH fringes were fringing.

The REFL path didn't need much, although we had never been going through the center of the HWP and PBS that are used to reduce the power before going to the PDs, so we translated them a millimeter or two.

We see signal on dataviewer for all of the channels that we're interested in....AS55 I&Q, ASDC, REFL11 I&Q, REFLDC (which comes from REFL55). 

Locking:

Locking MICH was very easy, after we rotated the phase of AS55 to get all the good MICH signal in the Q phase.  Part of the criteria for this was that the AS55_Q_ERR signal should cross zero when ASDC went to 0.  This was done very coarsely, so we need to do it properly, but it was enough to get us locked.  We changed the phase from 24.5 to 90 deg.

PRCL has been more of a challenge, although we're still working on it.

On the back face of the Faraday, we see the michelson fringes, but they are not getting through the Faraday's aperture.  This implies that we have a poorly aligned michelson, in that the interference between the returning beams from the ITMs is happening at a different place than the original beam splitting.  Yuta is working on getting a better MICH right now.  EDIT, 10 minutes later....   This seems to be fixed, and the MICH fringes enter the back aperture of the FI, but there is still the PRM refl problem (next paragraph).

Also, when we get the most bright REFL beam, we see that there is some very obvious clipping in the back of the Faraday aperture, and this is matched by a clipped-looking REFL beam on the AS table.  We must understand what we have done wrong, such that when the beam is actually going through the Faraday, we see a much dimmer beam.  It's possible that there is some clipping happening at that time with the in-vac REFL path....we need to check this.  It's not a clipping problem on the AS table - I checked, and the beams are still reasonably  well centered on all of the mirrors.

We think that the MICH / REFL beam problems may be that the input pointing is close, but not perfect.  We have not confirmed today that the beam is centered on ETMY.  We should do this as part of our final alignment procedure before putting on doors.

Plans for tomorrow:

Get POP aligned, especially the camera, so we can see what our intracavity mode really looks like in the PRC.  This is probably (in part, at least) due to our having moved PR2 around, so the transmitted beams aren't in exactly the same place.

We think that it's more useful in the short term to check out the PRC, and since the clipping problem with the REFL beam is likely an imperfect input pointing, we want to use the other measured G&H mirror, and do another half-PRC test, with the test mirror in front of the BS.  This requires much less perfection in the input pointing, so it should be very quick to set up.

Confirm that PRM oplev is still aligned (turn laser back on first).

Plans for next week:

Perfect the input pointing, by checking the beam position at ETMY.  Recheck all corner alignment.

Try again locking PRMI in air.  First, confirm ITM and BS oplevs are all aligned.

ELOG V3.1.3-